A rectangular field has an area of 600 square feet. If a scale drawing of the fied has a length of 4 inches and an area of 20 square inches, what is the perimeter of the field in real life?
Answer by Alan3354(69443)

A rectangular field has an area of 600 square feet. If a scale drawing of the fied has a length of 4 inches and an area of 20 square inches, what is the perimeter of the field in real life?
------------
Drawing is 4 by 5 inches.
---
The ratio of areas is 600*12*12/20 = 4320:1
Area is a function of L*W --> ratio of L & W is sqrt(4320) =~ 65.73:1
L = 4*65.73 inches = 21.91 feet
W = 5*65.73 inches = 27.386 feet